OpenStudy (anonymous):

The product of two consecutive even integers is 16 more then 8 times the smallest integer. Determine the integers.

Parth (parthkohli):

If x is the smaller, then x + 2 is the bigger one. Let's put up an equation here. \(\Large \color{MidnightBlue}{\Rightarrow x(x + 2) = 16 + 8x }\) \(\Large \color{MidnightBlue}{\Rightarrow x^2 + 2x = 16 + 8x }\) \(\Large \color{MidnightBlue}{\Rightarrow x^2 - 6x - 16 = 0 }\) Use the quadratic formula to solve.

Parth (parthkohli):

@mr.awesome two consecutive EVEN. I think you misread that part.

OpenStudy (anonymous):

\[n(n+2)=n+16 \] \[n^2+2n=n+16 \] \[n^2+n=16 \] \[n^2+n-16=0 \] \[x= \frac { -1\pm \sqrt { 65 }}{ 2}\\\\ x= \frac{ -1 + \sqrt{ 65 }} { 2 }\\\\ \]